OpenStudy (4everaddicted2anime):

Medal and Fan. Find the volume of a sphere that has a surface area of 16pi sq. in.

OpenStudy (anonymous):

surface area of sphere = 4 pi r^2 = 16 => r = 2/sqrt(pi) volume of sphere = 4/3 pi r^3 = 4/3 pi (2/sqrt(pi)^3 = 32/3sqrt(pi)